# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ Appendix D Additional Application Analyses
D.1 Whisker Activity
Since the authors were interested in how cells from this brain region encoded the sensory input of whisker activity, they recorded how much the whiskers moved when they were experimentally manipulated. The movement-by-moment activity of the whiskers can thus be used as a functional covariate to estimate how the association evolves across time. We fit the following concurrent model
$\text{logit}(\mathbb{P}[{Y}_{i,j,l}(s)\mid X_{i,j}(s)])=$ $~{}\beta_{0}(s)+X_{i,j}(s){\beta}_{1}(s),$
where $X_{i,j}(s)\in\mathbb{R}$ is a measurement of whisker activtiy at trial timepoint $s$ on trial $j$ for the animal that neuron $i$ was recorded from. We centered and scaled whisker activity for interpretability.
Figure 2: Whisker–Neural activity association. Functional coefficient estimates for Independent (Ind), Exchangeable (Exch), and Auto-regressive-1 (AR1) working correlation structures. Time relative to whisker stimulation.
Because $\widehat{{\beta}}_{1}(s)$ is significantly positive throughout the entire trial (i.e. across the functional domain), suggests that the neuron–whisker association is not driven by whisker stimulation. Interestingly the magnitude of association is roughly twice as high with the AR1 structure. Both the intercept and the functional slope suggest that the whisker stimulation (occurring at timepoint 00 sec) has only a small influence on the association between neural activity and whisker activity. Since $X_{i,j}(s)$ is centered and scaled, the functional intercept can be interpreted as the mean log odds of neural activity at an average speed. This appears to be affected by the stimulation as the shape of the functional intercept appears to rapidly increase and then decrease after stimulation (timepoint 00 sec). The AR1 analysis, for example, took $13.5$min, illustrating its speed for large datasets with functional covariates.

### v2

@@ Appendix D Additional Application Analyses
D.1 Background for Calcium Imaging Data Analysis
Since calcium imaging and electrophysiology record the activity of many neurons, and recordings are collected in several animals, analyses differ in how the target population is defined and the nesting of neurons within animal is modeled. For example, the neural pseudo-population strategy, as we refer to it, fits a single model to a dataset that pools neurons across animals (e.g. see Figures 1, 3, and 3 of Willmore et al. (2023); Zhang et al. (2023); Roesch et al. (2009), respectively). This conceptualizes neurons, both within and across animals, as exchangeable given covariates and model parameters. The animal-specific neural population strategy, as we refer to it, summarizes the collection of neurons separately in each animal, and then summarizes the animal-specific statistics with a secondary pooled test statistic (e.g. see Figs 2H in (Legaria et al., 2022), Fig 1G, 1I in Inácio et al. (2025)). The animal-level summary is usually a model fit to, or an average of, the activity of all neurons recorded from that animal (e.g. see Figs 2G in (Legaria et al., 2022)). This ignores uncertainty in the animal-level statistics when estimating a pooled test statistic. A third approach estimates a test statistic on data from each neuron separately and then fits a model to those statistics (e.g. see Figures 1K, 2E-H of Inácio et al. (2025)). The pooled test ignores uncertainty in the neuron-level statistics, and models the neuron-level statistics estimated on data from neurons in the same animal as independent.
Analysis strategies differ in how the longitudinal structure of experiments are modeled. One strategy is to treat the neural responses of cluster $i$ –– however defined –– as exchangeable across trials given model parameters (e.g. see Figure 3E of Jeong et al. (2022) for an example from photometry). A second strategy averages the response across trials and analyzes those trial-averaged measures (e.g. see Figure 2 of Coddington et al. (2023)). This discards longitudinal information. A third strategy accounts for the longitudinal structure with random effects Loewinger et al. (2025), yielding conditional estimates in binary outcome settings.
Analyses vary in how the densely-sampled neural time-series of each trial is conceptualized. Arguably, the most common strategy analyzes univariate summaries of the time-series (e.g. a trial firing rate for each neuron $i$ and trial $j$: $\bar{Y}_{i,j}=\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}Y_{i,j}(s)$) pooled across animals and/or trials (e.g. see Figures 1, 3, and 3 of Willmore et al. (2023); Zhang et al. (2023); Roesch et al. (2009), respectively). This strategy can obscure behavior–brain associations and substantially change scientific conclusions because it discards timing information about how covariate-outcome relationships evolve across trial timepoints (Loewinger et al., 2025). A second strategy is to retain the time-series structure, but model covariate-neural activity associations as constant across trial timepoints for each neuron. For example, Inácio et al. (2025) has the goal of identifying neurons associated with a particular behavior over time (e.g. see Figures 1-2). To that effect, they fit a Pearson correlation between behavior and neural activity in each cell separately. This is comparable to the linear regression $\mathbb{E}[\bm{Y}_{i,j}(s)\mid{x}_{i,j}(s)]=\gamma_{0}^{(i)}+\gamma_{1}^{(i)}x_{i,j}(s)$ (e.g. see figure 1K in Inácio et al. (2025)). This models the covariate–outcome relationship as constant across trial timepoints. A third strategy proposed in (Loewinger et al., 2025) addresses this by modeling each trial as a functional outcome. They demonstrate this strategy on fiber photometry data, a recording technique that yields a single neural signal per animal. It would be desirable to apply longitudinal FDA strategies to other recording techniques, such as calcium imaging or electrophysiology data. This has not been done, to our knowledge, and we believe this is primarily due to the fact that those modalities record potentially tens of thousands of signals per animal, and existing approaches may not scale.
D.2 Main Text Analyses with Different Correlation Structures
Here we show fGEE estimates fitted with a range of correlation structures to illustrate that the conclusions from the analysis in Figures 1(a)-1(b) are not sensitive to the working correlation structure.
(a) Speed–neural activity association.
(b) Whisker stimulation effect.
Figure S2: Functional coefficient one-step estimates with a working covariance $\mathbb{V}=\mathbb{V}_{Lon}\otimes\mathbb{V}_{Fun}$. “ar1 $\times$ ind” indicates that $\mathbb{V}_{Lon}$ has an AR1 structure, and $\mathbb{V}_{Fun}$ has a working independence correlation structure. (a) Fit times (sec) in order from top to bottom: 349.84, 394.15, 394.61, 188.86. (b) Fit times (sec) in order from top to bottom: 41.48, 36.64, 91.95, 16.65.